Gordon’s Smoothie Stand makes three types of smoothies: blueberry lemon, orange swirl, and triple berry. Before all flavors are added, the smoothies go through a joint mixing process that costs a total of $43 per batch. One batch produces 21.75 cups of blueberry lemon smoothies, 29 cups of orange swirl smoothies, and 36.25 cups of triple berry smoothies. In addition, Gordon has studiously noted that the mixing process necessary for triple berry and blueberry lemon smoothies takes twice as long as it does for orange swirl smoothies. Allocate the joint costs of production to each product using the weighted average method
Answer #1
Please give positive ratings so I can keep answering. It would help me a lot. Please comment if you have any query. Thanks!
Gordon’s Smoothie Stand
Allocation of joint costs
Blueberry lemon smoothies Orange swirl smoothies Triple berry smoothies Total Note
Number of cups produced                                          21.75                                  29.00                                  36.25 A
Weight                                             2.00                                    1.00                                    2.00 B
Weighted Number of cups produced                                          43.50                                  29.00                                  72.50               145.00 C=A*B
Cost per batch                 43.00 D
Cost per Weighted Number of cups produced                   0.30 E=D/C
Cost allocated to each product                                          12.90                                    8.60                                  21.50                 43.00 F=C*E